soving systems by substitution?
1. y=x+3
y=2x+4

Answer by checkley77(12569)

y=x+3
y=2x+4
x+3=2x=4
x-2x=4-3
-x=1
x=-1 ans.
y=-1=3
y=2 ans.
Proof:
2=2*-1+4
2=-2+4
2=2